### v1

# PhysLLM: Harnessing Large Language Models for Cross-Modal Remote Physiological Sensing

Abstract
Remote photoplethysmography (rPPG) enables non-contact physiological measurement but remains highly susceptible to illumination changes, motion artifacts, and limited temporal modeling. Large Language Models (LLMs) excel at capturing long-range dependencies, offering a potential solution but struggle with the continuous, noise-sensitive nature of rPPG signals due to their text-centric design. To bridge this gap, we introduce PhysLLM, a collaborative optimization framework that synergizes LLMs with domain-specific rPPG components. Specifically, the Text Prototype Guidance (TPG) strategy is proposed to establish cross-modal alignment by projecting hemodynamic features into LLM-interpretable semantic space, effectively bridging the representational gap between physiological signals and linguistic tokens. Besides, a novel Dual-Domain Stationary (DDS) Algorithm is proposed for resolving signal instability through adaptive time-frequency domain feature re-weighting. Finally, rPPG task-specific cues systematically inject physiological priors through physiological statistics, environmental contextual answering, and task description, leveraging cross-modal learning to integrate both visual and textual information, enabling dynamic adaptation to challenging scenarios like variable illumination and subject movements. Evaluation on four benchmark datasets, PhysLLM achieves state-of-the-art accuracy and robustness, demonstrating superior generalization across lighting variations and motion scenarios.
1 Introduction
Figure 1: The comparison between (a) typical CNN-based rPPG model, (b) pure LLM model, and (c) our proposed PhysLLM. And (d) is the performance comparison of different methods on the UBFC-rPPG [1] dataset.
Remote photoplethysmography (rPPG) is a non-contact technique that allows for the remote measurement of physiological signals such as heart rate (HR) [43, 3] and blood pressure [36] by analyzing subtle color changes in the skin caused by blood flow. Unlike traditional contact-based methods, such as electrocardiograms (ECG) and Photoplethysmograph (PPG) [20], rPPG does not require physical sensors attached to the body, making it a more convenient and less intrusive option for continuous health monitoring. Based on these advantages of rPPG, more and more excellent work and research have been committed to improving its accuracy, robustness and scalability in recent years.
Traditional remote photoplethysmography (rPPG) methods [23, 19, 4, 4, 34] typically rely on signal processing techniques to isolate rPPG signals from videos. More recent deep learning methods, including CNNs [38, 10, 17, 21, 26, 2, 16] and Transformers [42, 25, 24, 40], have been proposed to address some of these challenges by learning more complex representations of facial features. Fig. 1(a) shows a typical CNN-based framework where features extracted by the encoder are directly fed into the estimator to produce the final sequence. However, these methods remain vulnerable to visual noise (e.g., motion blur, occlusions, low-resolution inputs) and over-reliance on single video streams, limiting their robustness in practical scenarios. The exclusive dependence on visual data creates inherent performance boundaries, particularly in dynamic real-world environments. This underscores the necessity for advanced frameworks that integrate multi-modal inputs and adaptive feature learning through large-scale modeling.
Large Language Models (LLMs) offer significant advantages for enhancing rPPG methods, particularly in modeling long-term temporal dependencies [30, 8]. While traditional rPPG approaches struggle with extended video sequences, LLMs excel at capturing complex sequential patterns, making them promising candidates for physiological signal estimation. Previous works like TimeLLM [8] have demonstrated the effectiveness of LLMs in time-series applications, suggesting untapped potential for rPPG tasks. However, directly applying LLMs to rPPG estimation presents fundamental challenges, as illustrated in Fig. 1(b). The mismatch between LLMs’ discrete operations and rPPG features’ continuous nature leads to poor representations and high noise sensitivity. Despite LLMs’ proven capabilities in cross-modal tasks, their text-oriented architecture requires significant adaptation for effective physiological signal analysis.
To address these limitations, we propose PhysLLM, a collaborative optimization framework that integrates LLMs with specialized rPPG processing components. As shown in Fig. 1(c), to combine CNN’s local spatio-temporal feature extraction with LLM’s superior long-range temporal reasoning for more robust physiological measurements, we employ a CNN-based rPPG as the base model, leveraging both the multi-scale features and the estimated rPPG signals to guide LLM learning. Specifically, we introduce a Dual-Domain Stationary (DDS) Algorithm to stabilize the base model’s rPPG output and a Vision Aggregator module to fuse multi-scale hemodynamic features. To bridge the gap between rPPG features and LLM processing, we propose Text Prototype Guidance (TPG), which aligns sequence and multi-scale features with LLM text prototypes. Furthermore, we introduce task-specific cues, i.e., environmental factors, physiological knowledge, and task description, with learnable word vectors to enhance the LLM’s understanding of the rPPG context through cross-modal learning, integrating both visual and textual information. Therefore, PhysLLM adapts to varying conditions such as lighting changes and motion artifacts, significantly improving the accuracy and robustness of rPPG measurement. Overall, our contributions include:
  • We propose PhysLLM, the first framework integrating LLMs into rPPG measurement to establish interpretable connections between physiological dynamics and contextual semantics. The architecture’s inherent capacity for long-sequence dependency modeling enables superior performance in complex real-life scenarios.
  • We propose a novel time-frequency Dual-Domain stationary (DDS) Algorithm to address spectral-temporal instability through adaptive coefficient modulation with exponential decay characteristics. DDS ensures the processed time series to maintain periodic consistency while reducing noise interference.
  • We design task-specific cues to inject physiological priors through physiological statistics, environmental context and task description, enabling dynamic adaptation to challenging scenarios.
  • We propose the Text Prototype Guidance (TPG) strategy to establish cross-modal alignment by projecting hemodynamic features into LLM-interpretable semantic space, significantly reducing the gap between sequential, visual, and textual modalities.
  • Extensive experiments demonstrate the superiority of PhysLLM even under scenarios with serious degradation.
2 Related Work
Figure 2: Framework of the PhysLLM. The architecture of PhysLLM comprises three principal data streams that operate in concert. The leftmost stream represents the Physiological Cue-Aware Prompt Learning module, which incorporates task-specific prior knowledge through adaptive prompt learning while generating context-aware prompt tokens. The central and rightmost streams collectively form the Text-Vision-Sequence Embedding Generation pipeline, which integrates our novel Dual-Domain Stationary (DDS) Algorithm and Text Prototype Guidance (TPG) module. This integrated approach facilitates the extraction of sequence tokens from temporal physiological data and visual tokens from facial imagery, both guided by LLM-generated text prototypes that serve as semantic anchors for cross-modal alignment.
Remote Physiological Measurement. rPPG has advanced significantly from early signal processing methods to deep learning-based approaches. Traditional techniques, such as blind source separation and skin reflection-based models [23, 19, 4, 34], initially aimed to isolate physiological signals but struggled with real-world variability. Later methods incorporated physiological priors to improve robustness [45], yet they remained sensitive to noise and illumination. With the rise of deep learning, more recent approaches to rPPG measurement have leveraged the power of neural networks to capture rich spatio-temporal representations. Early deep learning-based methods introduced end-to-end spatiotemporal networks [26, 12, 38, 2, 21, 16, 40, 42]. These models demonstrated superior performance compared to traditional methods, as they were capable of capturing intricate temporal and spatial relationships within the video data. One major limitation is their sensitivity to visual interference, such as motion blur, occlusions, and varying lighting conditions, all of which can compromise the quality of the extracted rPPG signal. Recent advancements in vision-language multi-modal learning [44] have shown great promise in enhancing robustness by aligning with additional modalities for rPPG measurement.
LLMs for Time Series Tasks. Recent advancements in LLMs have demonstrated remarkable potential in time series analysis tasks, particularly through their ability to learn temporal patterns and perform cross-modal reasoning. Several pioneering works have explored reprogramming LLMs for time series forecasting without architectural modifications. For instance, Time-LLM [8] introduces a novel reprogramming framework that aligns time series embeddings with LLM token spaces using learnable prompt tokens, achieving state-of-the-art performance across multiple domains. This approach builds upon foundational insights from Time Series Forecasting with LLMs [6], which systematically evaluates LLMs’ inherent temporal reasoning capabilities and proposes specialized fine-tuning strategies to enhance their forecasting accuracy. In healthcare applications, LLMs have shown particular promise for physiological time series interpretation [14]. While existing research primarily focuses on standard physiological signals like ECG and EEG, the application of LLMs to rPPG analysis remains underexplored. Current approaches for rPPG-based health monitoring typically rely on specialized neural architectures, potentially overlooking the cross-modal generalization capabilities inherent in foundation models. Our work bridges this gap by adapting the time series reprogramming module while incorporating rPPG-aware vision-language knowledge, enabling LLMs to interpret subtle cardiovascular patterns from facial videos.
3 Methodology
The overall architecture of the proposed PhysLLM is designed to integrate textual, visual, and signal information into a unified framework, enabling robust rPPG physiological estimation. As illustrated in Fig. 2, the framework consists of two key components: Text-Vision-Sequence Embedding Generating (the right part of Fig. 2) and Physiological Cue-Aware Prompt Learning (the left part of Fig. 2). The detailed components of the model and the training objectives are described below.
3.1 Text-Vision-Sequence Embedding Generating
While current deep learning frameworks struggle to extract reliable rPPG signals from facial videos due to environmental interference, existing physiological priors in models like PhysNet [38] demonstrate inherent robustness to such variations. Building on this foundation, we establish a noval architecture with: 1) A fixed PhysNet [38] backbone preserving raw rPPG signals and spatio-temporal features, 2) Three trainable enhancement modules (Dual-Domain Stationary Algorithm module for stable signal tokenization, dedicated multi-scale interaction module for hierarchical feature extraction, Text Prototype Guidanc module aligning these representations with LLM semantic spaces) that progressively refine the physiological representations.
Dual-Domain Stationary (DDS) Algorithm. To effectively reduce the interference of noise on the model and further enhance its robustness and prediction accuracy, we propose a novel algorithm. This algorithm optimizes the signal processing workflow in a targeted manner, significantly improving the quality of rPPG signals and providing a more reliable data foundation for subsequent analysis.
Specifically, let $x\in\mathbb{R}^{B\times L}$ denote the raw rPPG waveform extracted by the PhysNet backbone [38], where batch size $B$ and sequence length $L$ inherit the output dimensions of the backbone. We first compute the global mean $\mu$ and global standard deviation $\sigma$ across the entire sequence,
The normalized output $x^{\prime}$ is then calculated as:
$\small x^{\prime}=\frac{x-\mu}{\sigma+\epsilon},\quad\epsilon=10^{-5},\vspace{-0.3em}$ (1)
We denote similar standardization processes in a unified manner as function $\mathcal{S}(\cdot)$. To ensure global stationarity, we proceed temporal smoothing with the following operations:
$\small z^{time}_{i}=\alpha\cdot x^{\prime}_{i}+(1-\alpha)\cdot z_{i-1},\quad z_{0}=x^{\prime}_{0},\vspace{-0.2em}$ (2)
where i denotes the current frame value, and $\alpha$ is a smoothing factor. The proof of stationarity is provided in the Appendix. In the following text, we denote these smoothing operations collectively as $\mathcal{Z}(\cdot)$. Therefore, the above operation can be simplified as $z^{time}=\mathcal{Z}(x^{\prime})$.
In parallel, the module performs frequency domain decomposition using the discrete wavelet transform (DWT). The DWT decomposes the input signal into approximation coefficients ($ac$) and detail coefficients ($dc$) on multiple scales. Specifically, the DWT can be expressed as:
$\small x_{ac},[x_{dc,1},\dots,x_{dc,J}]=\text{DWT}(x),\vspace{-0.2em}$ (3)
where $J$ is the decomposition level ($J=3$ in this implementation). After normalization, the inverse wavelet transform (IDWT) reconstructs the smoothed frequency-domain representation $z^{fre}$:
$\small z^{fre}=\text{IDWT}(\mathcal{Z}(\mathcal{S}(x_{ac})),[\mathcal{Z}(\mathcal{S}(x_{dc,1})),\dots,\mathcal{Z}(\mathcal{S}(x_{dc,J}))]).\vspace{-0.2em}$ (4)
To combine the advantages of time-domain and frequency-domain processing, an adaptive weighting mechanism is introduced. The final smoothed output $x_{rec}$ is computed as:
$\small z=(1-\text{$\beta$})\cdot z^{time}+\text{$\beta$}\cdot z^{fre},\vspace{-0.2em}$ (5)
[…]
Figure 3: Architecture of the Vision Aggregator. It employs a hierarchical attention architecture to dynamically synthesize multi-scale feature representations.
Multi-scale Interaction via Vision Aggregator (VA). We introduce the multi-scale Interaction module to effectively integrate multi-scale features from different modalities. As shown in Fig. 3, this module leverages a combination of cross-attention and self-attention mechanisms to capture both inter-modal and intra-modal relationships between high-level and low-level features. By employing learnable scaling parameters, the module ensures that the fused feature representation is adaptive and context-aware.
Specifically, we extract $M$ layers features from the backbone, denoted as $\mathcal{F}=[f_{1},f_{2},f_{3},...,f_{M}]$, in which $f_{i}\in\mathbb{R}^{B\times T\times H\times W}$. To align these features into a common embedding space, we employ a set of feature projection layers, ${F_{i}}=\text{Projection}(f_{i},l_{target})$, where $l_{target}$ is the setting length after compression.
Since the strong semantic alignment between deep visual features and text space, we use the deep feature $F_{M}$ as queries to dynamically extract missing details from shallow features $X=\mathrm{Concat}(F_{1},F_{2},\dots,F_{M-1})$. This results in a visual feature $F_{cross}\in R^{B\times T\times D}$ with richer fine-grained features. This process can be formulated as:
[…]
where CrossAttention(·) denotes the cross-attention mechanism, Concat(·) represents the concatenation operation.
To enhance the representation by capturing internal dependencies within the cross-attended feature, we incorporate a self-attention mechanism into the feature map $F_{self}$, formulated as:
[…]
To combine the outputs of the cross- and self-attention mechanisms, the final fused features are computed as:
[…]
where $\gamma_{1}$, $\gamma_{2}$ are the learnable vector.
Text Prototype Guidance (TPG). To fully leverage the prior knowledge embedded in rPPG signals and visual features, we aim to enable these modalities to play a critical role in the task. However, rPPG signals and visual features are neither directly editable nor easily describable in natural language without missing information, which poses significant challenges for guiding LLMs to understand temporal and visual features without resource-intensive fine-tuning. Specifically, the non-discrete and high-dimensional nature of such data makes it difficult to transform them into symbolic representations suitable for language models.
As shown in Fig. 2(c), to bridge this gap, we propose to reprogram the rPPG signals and visual features with word embeddings ${E}\in\mathbb{R}^{V\times D}$, where $V$ is the vocabulary size. Nevertheless, there is no prior knowledge indicating which source tokens are directly relevant. Thus, simply leveraging $E$ will resulting large and potentially dense reprogramming space. A simple solution is to maintain a small collection of text prototypes by linearly probing, denoted as ${E^{{}^{\prime}}}\in\mathbb{R}^{V^{{}^{\prime}}\times D}$, where $V^{\prime}\ll V$.
The text prototypes ${E^{{}^{\prime}}}$ are then used to enhance the interaction between rPPG signals, visual features, and language model. To extract more sufficient details from the input video and signal, we devise a block, which consists of multiple different transformer layers. Specifically, given input feature $X$ :
$\small X_{self}=\text{SelfAttention}(\text{$X$}),\vspace{-0.2em}$ (9)
$\small{E^{{}^{\prime}}_{fusion}}=\text{${E^{{}^{\prime}}}+X_{self}$},\vspace{-0.2em}$ (10)
[…]
$\small{Output}=\text{${E^{{}^{\prime}}_{fusion}}$ + $y_{2}(E^{{}^{\prime}}_{fusion};X)$},\vspace{-0.2em}$ (12)
$\small\mathcal{T}_{out}=\text{$FFN({E^{{}^{\prime}}_{2}})$},\vspace{-0.2em}$ (13)
where ${E^{{}^{\prime}}_{1}}$ and $E^{{}^{\prime}}_{2}$ are the updated versions of $E^{{}^{\prime}}$. It is worth noting that the TPG module guides not only visual features, but also sequence features, so the input feature X here has the both two meanings. Also, the part of visual features and the part of sequence features shared parameters to learn potential associations.
For simplicity, we represent this sequence of operations as a single function $TPG(\cdot)$, so the output of this module can be expressed as $\mathcal{T}_{out}=TPG(X)$.
Figure 4: Introduction to the composition of cues. (a) Extracting visual priors (e.g., lighting, facial expressions, occlusions) via LLaVA and encoding them into visual tokens. (b) Tokenizing textual descriptions of the rPPG task to derive task-specific priors. (c) Analyzing statistical features of rPPG signals from the backbone network to generate statistical prior tokens. These cues synthesize visual, semantic, and statistical priors for enhanced physiological signal analysis.
Table 1: Intra-dataset testing results on UBFC-RPPG [1], PURE [27], BUAA [37] and MMPD [31] datasets. Best results are marked in bold.
Method UBFC-rPPG PURE BUAA MMPD
[…]
ICA [22] 16.00 25.65 0.44 4.77 16.07 0.72 - - - 18.60 24.30 0.01
CHROM [5] 4.06 8.83 0.89 5.77 14.93 0.81 - - - 13.66 18.76 0.08
Green [33] 19.73 31.00 0.37 10.09 23.85 0.34 6.89 10.39 0.60 21.68 27.69 -0.01
POS [35] 4.08 7.72 0.92 3.67 11.82 0.88 - - - 12.36 17.71 0.18
Meta-rPPG [9] 5.97 7.42 0.57 2.52 4.63 0.98 - - - - - -
Contrast-Phys+ [29] 0.21 0.80 0.99 0.48 0.98 0.99 - - - - - -
PhysNet [39] 2.95 3.67 0.97 2.10 2.60 0.99 10.89 11.70 -0.04 4.80 11.80 0.60
PhysFormer [41] 0.92 2.46 0.99 1.10 1.75 0.99 - - - 11.99 18.41 0.18
Efficientphys [13] 1.41 1.81 0.99 4.75 9.39 0.99 16.09 16.80 0.14 13.47 21.32 0.21
PhysLLM (Ours) 0.21 0.57 0.99 0.17 0.35 0.99 6.48 8.48 0.63 4.36 10.76 0.65
3.2 Physiological Cue-Aware Prompt Learning
The extraction of reliable rPPG signals from facial video data presents significant challenges due to varying lighting, subject mobility, and diverse skin tones in complex real-world scenarios. To address these challenges through cross-modal learning, we propose a Physiological Cue-Aware Prompt Learning framework that generates rPPG sample-specific cue and make them guide adaptive token learning.
Cue Caption. Prompt-based adaptation enables LLMs to address specialized subtasks without requiring parameter updates, as demonstrated by the zero-shot capabilities of architectures like LLaMA [32] and DeepSeek [7]. To eliminate manual annotation of context-aware descriptors, we leverage LLaVA [11] for automated physiological cue generation through structured prompt engineering, as shown in Fig. 2(b). For rPPG-specific adaptation, our prompts focus on three critical visual domains: 1) facial anatomical characteristics, 2) transient emotional expressions, and 3) environmental illumination dynamics, as exemplified by our designed query templates in Fig. 4.
Formally, given an input video sequence $V\in\mathbb{R}^{3\times T\times H\times W}$, we extract the central frame $I_{t}$ and process it through LLaVA with task-oriented prompts to obtain detail description, and tokenize the description,
$\small\mathcal{C}_{vision}=Tokenizer(LLaVA(I_{t},Q)),\vspace{-0.2em}$ (14)
where $Q$ denotes our physiological cue queries.
Furthermore, we formalize domain knowledge through task-specific primers derived from consensus descriptions in rPPG literature. These primers are encoded into LLM-compatible tokens via the model’s native tokenizer:
$\small\mathcal{C}_{task}=Tokenizer(\mathcal{P}),\vspace{-0.2em}$ (15)
where $\mathcal{P}$ represents the standardized task preamble.
While traditional text-based prompting strategies have demonstrated effectiveness in general vision-language tasks, we identify critical limitations when applied to physiological signal estimation: (1) Static visual descriptors cannot sufficiently capture temporal hemodynamic variations, and (2) Task-agnostic queries fail to address rPPG-specific contextual dependencies. To bridge this semantic gap, we propose statistical priors derived from pretrained rPPG models as physiological knowledge supplements. Although ground truth signals remain unavailable during inference, the pretrained network’s output sequences (discussed in Section 3.1) preserve valid distribution characteristics. So the statistical cue $\mathcal{S}$ is derived from the pretrained model’s rPPG signal projections $x_{\text{enc}}\in\mathbb{R}^{B\times T}$, where $B$ denotes batch size and $T$ the temporal dimension. For each sample $b\in[1,B]$, we compute,
[…]
$\Gamma(x_{\mathrm{enc}}[b,:]),\ \mathrm{sgn}\left(\sum_{t=1}^{T-1}(x_{\mathrm{enc}}[b,t+1]-x_{\mathrm{enc}}[b,t])\right),$
$\mathrm{TopK}\left(\mathcal{L}(x_{\mathrm{enc}}[b,:]),5\right)\Biggr{\}}$
where $\Gamma(\cdot)$ calculates signal trends through first-order differencing,
$\small\Gamma(\mathbf{x})=\sum_{i=2}^{T}(\mathbf{x}_{i}-\mathbf{x}_{i-1}).\vspace{-0.2em}$ (17)
Then $\mathcal{S}$ will make up the symbolic representation as described in Fig. 4(c). Finally, the tokenization process is formulated as $\mathcal{C}_{static}=\mathrm{Tokenizer}\left(\mathcal{S}\right).$
Adaptive Prompt Learning (APL). Conventional multimodal fusion approaches typically employ static weighting coefficients for prompt integration, suffering from two limitations: (1) fixed combination ratios cannot adapt to learnable LLM, and (2) inability to adaptively select the key cues across different data. As shown in Fig. 2(a), our adaptive fusion paradigm addresses these constraints through learnable hierarchical composition. Formally, having obtained three fundamental cue captions $\mathcal{C}=\{\mathcal{C}_{task},\mathcal{C}_{vision},\mathcal{C}_{static}\}$, we first implement modality-specific conditioning via dedicated transformation networks,
[…]
where each compressor employs temporal-contextual attention to maintain modality-specific patterns,
$\small\operatorname{AttentiveCompressor}_{k}(x)=\operatorname{Softmax}\left(\frac{Q_{k}K_{k}^{T}}{\sqrt{d}}\right)V_{k},\vspace{-0.3em}$ (19)
with ${Q_{k},K_{k},V_{k}}$ derived from $x$ through linear projections.
In addition, the adaptive fusion mechanism learns three independent parameter matrices:
$\small\mathbf{W}=\left[W_{\text{task}},\ W_{\text{vision}},\ W_{\text{static}}\right]\in\mathbb{R}^{3\times B\times L\times d},\vspace{-0.3em}$ (20)
where $L$ is predefined hyperparameter about target prompt sequence length, $d$ is the token embedding dimension, and $W_{\text{task}},W_{\text{vision}},W_{\text{static}}$ denote learnable tokens, respectively.
The complete fusion operation is formalized as:
$\small\mathcal{T}_{\text{cue }}=\sum_{\begin{subarray}{c}k\in\Omega\end{subarray}}\mathbf{W}^{(k)}\odot\mathcal{E}_{k},\Omega\triangleq\{task,\\ vision,\\ static\}.\vspace{-0.3em}$ (21)
3.3 Training Objectives
In Section 3.1, we obtain the stationary signal $z$ and the fused multi-scale visual features $F_{visual}$ through the DDS module and multi-scale interaction, respectively. Through TPG, we can obtain the text-guided vision token $\mathcal{T}_{\text{vision }}=TPG(F_{visual})$, and the text-guided signal token $\mathcal{T}_{\text{signal }}=TPG(z)$. Then, we input $\mathcal{T}_{\text{cue }}$, $\mathcal{T}_{\text{vision }}$, $\mathcal{T}_{\text{signal }}$ into LLM and predict waveforms $\hat{y}$. Finally, we employ the mean square error loss function as the total loss,
$\small\mathcal{L}_{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\vspace{-0.3em}$ (22)
where $y$ is the ground truth PPG signals.
4 Experimental Results
4.1 Datasets and Performance Metrics
We comprehensively assess the performance of PhysLLM on four widely used benchmark datasets (UBFC-rPPG [1], PURE [27], BUAA [37], and MMPD [31]), in comparison with state-of-the-art methods. Following [28, 38], we calculate mean absolute error (MAE), root mean square error (RMSE), and Pearson’s correlation coefficient (R) between the predicted HRs versus the ground-truth HRs as evaluation metrics. Notably, for MAE and RMSE, lower values indicate reduced error margins, whereas for R, values approaching 1.0 signify diminished error. Among them, both MAE and RMSE are measured in terms of bpm (beats per minute). And for convenience, these units will not be repeated in the subsequent tables or analyses.
4.2 Implementation Details
We conduct experiments on Pytorch and mainly based on the open-source toolkit rPPG-Toolbox [15]. For data pre-processing, we crop the face region in the first frame for each video clip and fix the region box in the following frames. Subsequently, we randomly sample a video chunk of 128 frames and resize them into 128 × 128 pixels. We use the default hyperparameters settings $\alpha$ = 0.8 and $l_{target}$ = 32. The backbone of our PhysLLM is based on the pre-trained PhysNet [38], where the training strategies follow the methodology outlined in the original paper. In the intra-dataset testing and ablation study, we use the same training data for PhysNet as for PhysLLM, ensuring consistency in the data used. For cross-domain experiments, we use the pre-trained PhysNet model trained on the PURE dataset [27], ensuring no test set overlap and consistent pre-training conditions. Furthermore, we use the Deepseek-1.5B [7] version as the LLM. The PhysLLM is trained with Adam optimizer and the initial learning rate and weight decay are 1e-4 and 5e-5, respectively. We train our model for 20 epochs on a NVIDIA A100 GPU with batch size of 4.
Table 2: Cross-domain generalization evaluation with dual-source training and single-target testing.
Method P+B $\to$ M P+U $\to$ M B+U $\to$ M
[…]
Efficientphys [13] 11.9 18.5 11.8 18.9 15.5 20.8
PhysFormer [41] 13.9 18.6 11.4 17.5 13.2 16.5
PhysNet [39] 13.2 16.7 11.0 17.3 13.5 17.0
PhysLLM (Ours) 11.9 15.3 9.95 14.96 12.1 15.2
Table 3: Cross-domain generalization evaluation with three-source training and single-target testing.
Method Others $\to$ MMPD Others $\to$ BUAA
[…]
Efficientphys [13] 13.2 20.02 0.08 32.3 34.0 -0.03
PhysFormer [41] 13.9 19.3 0.06 7.65 12.4 0.25
PhysNet [38] 12.8 16.3 0.12 12.75 16.37 0.08
PhysLLM (Ours) 12.2 15.5 0.19 6.0 8.6 0.63
4.3 Intra-dataset Testing
We first evaluate the HR estimation on all datasets under intra-dataset setting. We compare our method with 10 methods, including traditional methods, Green [33], ICA [22], CHROM [5], POS [35], and deep learning methods, Meta-rPPG [9], Contrast-Phys+ [29], PhysNet [39], PhysFormer [41], Efficientphys [13]. Table 1 displays intra-dataset testing results for UBFC-rPPG [1], PURE [27], BUAA [37] and MMPD [31].
HR Estimation on UBFC-rPPG [1]. On the UBFC-rPPG [1] dataset, we use the videos of the first 30 subjects for training and the videos of the remaining 12 subjects for testing according to  [18]. The HR estimation results are shown in Table 1. The proposed PhysLLM outperforms the state-of-the-art methods, and achieves MAE (0.21 bpm), RMSE (0.57 bpm), and R (0.99) metrics for HR prediction. The results show that PhysLLM not only fits the rPPG signal consistently but also tracks HR changes effectively over a long period.
HR Estimation on PURE [27]. Following the protocol in  [18], we compare PhysLLM with 9 methods. It can be seen from Table 1 that PhysLLM achieves the best performance on all evaluation metrics for HR, especially compared with the second-best method PhysFormer [41]. PhysLLM achieves improvements of 0.31 bpm in MAE and 0.63 bpm in RMSE, indicating the excellent robustness of PhysLLM against the interference induced by head movements.
HR Estimation on BUAA [37]. It is partitioned in sequence into training and test sets with a ratio of 7:3. The performance of existing methods on this dataset is reimplemented by ourselves with the rPPG toolbox [15]. The HR estimation results are shown in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (6.48 bpm), RMSE (8.48 bpm), and R (0.63) metrics for HR prediction. The results show the strong robustness of PhysLLM against various lighting condition disturbances.
HR Estimation on MMPD [31]. Following the protocol in  [46], we compared our method against existing state-of-the-art approaches, as depicted in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (4.36 bpm), RMSE (10.76 bpm), and R (0.65) metrics for HR prediction. These results demonstrate that PhysLLM performs well under real-world conditions.
4.4 Cross-dataset Testing
We also perform cross-dataset testing across UBFC-rPPG [1], PURE [27], BUAA [37], and MMPD [31] datasets to assess the generalization capability of PhysLLM. We conduct both Two-training and One-testing protocol and Three-training and One-testing protocol. In each case, we train on relatively simple datasets and evaluate performance on one more complex dataset.
Table 4: Ablation results of the main components on UBFC-rPPG [1].
Method PURE
[…]
$\times$ $\times$ 0.36 1.12 0.98
$\times$ $\times$ 0.41 1.26 0.98
$\times$ $\times$ 0.32 1.00 0.98
$\times$ 0.27 0.92 0.98
$\times$ 0.34 1.05 0.99
$\times$ 0.25 0.76 0.99
0.21 0.57 0.99
Dual-source training and single-target testing protocol.  To explore the generalization capacity of PhysLLM, we train on two datasets and uniformly test on the most difficult dataset, MMPD [31]. For example, $P+U\to M$ means training on PURE [27] and BUAA [37] datasets and testing on MMPD [31], and both the training and testing take the entire dataset. The two-training and one-testing dataset HR estimation results of our method and baseline methods are presented in Table 2. Despite more difficult cross-domain settings, PhysLLM still achieves the best performance, even reducing the MAE by 1.05 bpm and the RMSE by 2.34 bpm on $P+U\to M$, reflecting the superior generalization capacity of PhysLLM even on complex unseen domains, and the effectiveness of the visual contexts and LLM modeling.
Three-source training and single-target testing protocol. To further evaluate whether PhysLLM can learn domain invariance knowledge without mixing domain-specific knowledge, we trained on three datasets and tested on the remaining one dataset. For example, $Others\to MMPD$ means training on the other datasets (UBFC-rPPG [1], PURE [27], BUAA [37]) and test on MMPD [31]. It is worth noting that we only tested two difficult datasets: BUAA [37] and MMPD [31]. It can be seen from the results in Table 3 that PhysLLM achieves the best performance in both datasets, and even outperforms the existing state-of-the-art methods on MAE (6.0 bpm), RMSE (8.6 bpm), and R (0.63) metrics for HR prediction on $Others\to BUAA$. This shows that PhysLLM will not confuse domain-specific knowledge even if trained on three datasets, and can stably learn domain-invariant rPPG knowledge.
Table 5: Ablation study of prompt components on UBFC-rPPG [1].
Method UBFC-rPPG
[…]
$\times$ $\times$ 0.83 1.53
$\times$ $\times$ 0.71 2.26
$\times$ $\times$ 0.84 2.41
$\times$ 0.73 2.29
$\times$ 0.43 1.20
$\times$ 0.61 1.53
$\times$ 1.31 2.10
0.21 0.57
4.5 Ablation Study
Impact of Dual-Domain Stationary Algorithm (DDS). The role of DDS is to transform the input signal $x$ into a stationary signal $Z(x)$. To validate the effectiveness of the DDS, we construct a comparative experiment designed to remove DDS. Specifically, we directly use the rPPG signal from the video encoder, instead of the rPPG signal smoothed by DDS. The results are shown in the sixth row of Table  4. DDS reduces the MAE by 0.04 bpm, and the RMSE by 0.19 bpm on PURE dataset.
Impact of Vision Aggregator (VA). In this ablation study, we simply remove the Vision Aggregator to evaluate its impact on the model. As shown in the results on UBFC-rPPG dataset, the absence of the Vision Aggregator leads to a noticeable decrease in performance, underscoring its importance in the model architecture. The comparison between the fifth and seventh rows in Table 4 clearly demonstrates that the fusion of visual features is effective.
Impact of Text Prototype Guidance (TPG). To evaluate the effectiveness of text prototype guidance, we conducted an ablation study by removing the text prototype guidance. The results on UBFC-rPPG dataset are shown in the fourth row of Table  4. TPG brings a significant decrease in MAE (0.06 bpm) and RMSE (0.35 bpm). These results underscore the importance of SFD in facilitating the fusion of quasi-periodic video features.
Impact of prompt components. To evaluate the influence of different prompt components (Vision, Statics, Task, and adaptive learning) in cue caption and adaptive prompt learning of PhysLLM, we conducted ablation on them and presented in Table. 5 using MAE and RMSE metrics. The results reveal that the optimal configuration includes all the components achieving remarkable performance with MAE 0.21 bpm and RMSE 0.57 bpm. This represents a substantial 76% improvement over the worst-performing combination (MAE 1.31 bpm, RMSE 2.10 bpm), which utilized only Vision, Task and Statics components while excluding adaptive learning. This shows that our adaptive learning part is essential and critical. Several significant patterns emerge: configurations incorporating the Task component generally outperform those without it, suggesting the critical importance of task-specific prompting; the combination of Vision and Task is higher than that of Statics and Task, suggesting the generation of visual prompt is more in line with the relevant rules of rPPG tasks. Moderate performance (0.71-0.84 MAE) is achieved with Statics and adaptive learning only, but none approach the effectiveness of the optimal Task and adaptive learning pairing. These findings demonstrate the crucial importance of careful prompt component selection and emphasize that task-specific adaptability should be prioritized in prompt design, while also emphasizing the importance of learning headers assigning different prompt weights to different datasets and scenarios.
Figure 5: Visualization of saliency maps from PhysLLM on PURE [27] and UBFC-rPPG [1].
4.6 Visualization and Discussion
Following [28], the saliency maps of the PhysLLM are visualized in Fig. 5 using samples from UBFC-rPPG [1] and PURE [27] datasets. The more obvious the blue-green area, the more attention it means. We selected images with certain features from the dataset to test rPPG tasks such as head rotation, with a noticeable beard, darker skin tone, and wearing glasses. From the test results, it can be seen that physLLM can effectively capture rPPG-related parts. For example, the observable parts are concentrated on the cheeks and forehead, which is consistent with the relevant prior knowledge of rPPG. At the same time, relevant parts can effectively avoid hair and obstruction. For example, there is no prominent attention in the area where the hair appears, and even subtle changes in the neck can be observed. This is enough to reflect the robustness and effectiveness of PhysLLM in different scenarios.
5 Conclusions and Future Work
We propose PhysLLM, a framework that integrates LLMs with specialized rPPG components. The Text Prototype Guidance strategy bridges the cross-modal gap, while the Dual-Domain Stationary Algorithm addresses signal instability. Task-specific priors enhance adaptability in challenging scenarios. Experimental results across four datasets show PhysLLM achieves superior accuracy and robustness, particularly under variable illumination and motion. Future work includes improving cross-modal alignment and developing lightweight models for edge deployment.